We keep a switch at point X for a sufficient time and there after connect it to Y. In the circuit shown below . What amount of current will be measured through the resistor after 5. 0 micro second from t not if V = 9V R= 10 and C = 1 micro Farads?
1
Expert's answer
2021-08-11T14:47:22-0400

"I_0=\frac{V}{R}=\frac{9}{10}=0.9A"

"I=I_0e^\frac{-t}{Rc}"

Put value

"I=0.9e^\frac{-5\times10^{-6}}{10\times10^{-6}}"

"I=0.9e^{-0.5}"

"I=0.9\times0.606=0.5458A"